# MS-TVNet:A Long-Term Time Series Prediction Method Based on Multi-Scale Dynamic Convolution

### v1

@@ Abstract
Long-term time series prediction has predominantly relied on Transformer and MLP models, while the potential of convolutional networks in this domain remains underexplored. To address this gap, we introduce a novel multi-scale time series reshape module, which effectively captures the relationships among multi-period patches and variable dependencies. Building upon this module, we propose MS-TVNet, a multi-scale 3D dynamic convolutional neural network. Through comprehensive evaluations on diverse datasets, MS-TVNet demonstrates superior performance compared to baseline models, achieving state-of-the-art (SOTA) results in long-term time series prediction. Our findings highlight the effectiveness of leveraging convolutional networks for capturing complex temporal patterns, suggesting a promising direction for future research in this field.The code is realsed on https://github.com/Curyyfaust/TVNet.
@@
Long-Term Time Series Forecasting, CNN, Multi-Scale, Fourier transform
@@ I Introduction
Long-term time series prediction is a crucial task with extremely wide applications in professional fields such as energy management[1][2], weather forecasting[3], and health monitoring[4]. However, long-term time series forecasting tasks often involve complex multi-period scales and cross-variable dependencies[5], making the development of effective forecasting models a significant challenge. Real-world time series exhibit diverse variations across different temporal scales. To perform accurate time series forecasting, it is essential to adopt multi-scale modeling approaches that can capture features and dependencies from various time intervals. As illustrated in Figure 1, the same time series can be divided into smaller segments (in blue) or larger segments (in yellow), resulting in fine-grained or coarse-grained temporal features. In the figure, black arrows represent relationships between adjacent time steps, capturing local details, while colored arrows indicate relationships spanning longer time steps, reflecting global dependencies[6][7][8].

### v2

@@ Abstract
Long-term time series prediction has predominantly relied on Transformer and MLP models, while the potential of convolutional networks in this domain remains underexplored. To address this gap, we introduce a novel multi-scale time series reshape module, which effectively captures the relationships among multi-period patches and variable dependencies. Building upon this module, we propose MS-TVNet(Multi-Scale Time-Video Network), a multi-scale 3D dynamic convolutional neural network. Through comprehensive evaluations on diverse datasets, MS-TVNet demonstrates superior performance compared to baseline models, achieving state-of-the-art (SOTA) results in long-term time series prediction. Our findings highlight the effectiveness of leveraging convolutional networks for capturing complex temporal patterns, suggesting a promising direction for future research in this field.The code is released on https://github.com/Curyyfaust/TVNet.
@@
Long-Term Time Series Forecasting, CNN, Multi-Scale, Fast Fourier Transform;Dynamic Convolution
@@ I Introduction
Long-term time series prediction is a crucial task with extremely wide applications in professional fields such as energy management[1][2], weather forecasting[3], and health monitoring[4]. However, long-term time series forecasting tasks often involve complex multi-period scales[5][6] and cross-variable dependencies[7], making the development of effective forecasting models a significant challenge. Real-world time series exhibit diverse variations across different temporal scales. To perform accurate time series forecasting, it is essential to adopt multi-scale modeling approaches that can capture features and dependencies from various time intervals. As illustrated in Figure 1, the same time series can be divided into smaller segments (in blue) or larger segments (in yellow), resulting in fine-grained or coarse-grained temporal features. In the figure, black arrows represent relationships between adjacent time steps, capturing local details, while colored arrows indicate relationships spanning longer time steps, reflecting global dependencies[8][9][10].